OpenStudy (anonymous):

solve the equation 1/2x=3

OpenStudy (anonymous):

x=1.5 or 1 1/2 however you want to put it

OpenStudy (anonymous):

Yeah, it would be considered the division property of equality, so divide both sides by 1/2

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

thank you !!!!!!

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):



OpenStudy (anonymous):

lol, nice

OpenStudy (anonymous):

Just divided 3 on 1/2 caz : Look at this examle : we know : 2*2=4 So we we say : 2x=4 x=2 caz we can say : x=4/2 x=2 K ?:)

OpenStudy (anonymous):

good example